Hi,
I have a quadratic factorization question that I don't know how to do...
It's a Year 11 Maths question:
4x^4-37x^2+9
I have to factorize that and I would show you what I've done so far but I have no idea how to do this question...
Thank you so much for your help! :)

Answer by mananth(16946)

Hi,
I have a quadratic factorization question that I don't know how to do...
It's a Year 11 Maths question:
4x^4-37x^2+9
I have to factorize that and I would show you what I've done so far but I have no idea how to do this question...
Thank you so much for your help
4x^2-37x^2+9=0
4x^4-36^2 -x^2+9=0
4x^2(x^2-9)-1(x^2-9)=0
(4x^2-1)(x^2-9)=0
(2x+1)(2x-1)(x+3)(x-3)=0